OpenStudy (anonymous):

Solve. x2 + 10x = -2

OpenStudy (anonymous):

you want to use the formula again, or you want to complete the square? since \(10\) is even it is actually quicker to complete the square i can walk you through it if you like, or you can use the formula

OpenStudy (anonymous):

probably completely the square /:

OpenStudy (anonymous):

k takes just a couple steps 1) what is half of 10?

OpenStudy (anonymous):

that would definitely be 5

OpenStudy (anonymous):

right (math skillz) 2) what is \(5^2\) ?

OpenStudy (anonymous):

now that would definitely be 25

OpenStudy (anonymous):

right again we go from \[ x2 + 10x = -2\] directly to \[(x+5)^2=-2+25\\ (x+5)^2=23\]

OpenStudy (anonymous):

and thats -5 +/- sqrt23 right?

OpenStudy (anonymous):

two steps remain 3) take the square root of both sides (don't forget the \(\pm\) \[x+5=\pm\sqrt{23}\]

OpenStudy (anonymous):

yeah what you said snappy right?

OpenStudy (anonymous):

super snappy

OpenStudy (anonymous):

you want to make me happy you gotta make is snappy (george clinton)